# Boosting Transformer’s Robustness in PPG Signals Artifact Detection with Self-Supervised Learning

### v1

@@ Abstract
Recent research at CHU Sainte-Justine’s Pediatric Critical Care Unit (PICU) has revealed that traditional machine learning methods, such as semi-supervised label propagation and K-nearest neighbors, outperform Transformer-based models in artifact detection from PPG signals, mainly when data is limited. This study addresses the underutilization of abundant unlabeled data by employing self-supervised learning (SSL) to extract latent features from these data, followed by fine-tuning on labeled data. Our experiments demonstrate that SSL significantly enhances the Transformer model’s ability to learn representations, thereby improving its robustness in artifact classification tasks. Among various SSL techniques—including masking, contrastive learning, and DINO (self-distillation with no labels)—contrastive learning exhibited the most stable and superior performance in small PPG datasets. Further, we delve into optimizing contrastive loss functions, which are crucial for contrastive SSL. Inspired by InfoNCE, we introduce a novel contrastive loss function that facilitates smoother training and better convergence, thereby enhancing performance in artifact classification. In summary, this study establishes the efficacy of SSL in leveraging unlabeled data, particularly in enhancing the capabilities of the Transformer model. This approach holds promise for broader applications in PICU environments, where annotated data is often limited.
@@ I Introduction
Recently, the Pediatric Intensive Care Unit (PICU) at CHU Sainte-Justine (CHUSJ) has made notable advancements by developing a high-resolution research database (HRDB) [1, 2]. This innovative database directly integrates biomedical signals from various monitoring devices into the electronic patient record, ensuring seamless data incorporation throughout a patient’s PICU stay [3]. The integration of HRDB has significantly enhanced the Clinical Decision Support System (CDSS) at CHUSJ, boosting patient safety and underpinning decision-making with robust evidence [4]. In this context, early and accurate diagnosis of acute respiratory distress syndrome (ARDS) is a pivotal goal of the CDSS at CHUSJ. Oxygen saturation (SpO2) values, critical in ARDS diagnosis, are central to predicting and managing ARDS [5, 6], and play a key role in determining respiratory support strategies [7, 8, 9]. Moreover, the potential to predict SpO2 from Photoplethysmography (PPG) waveforms and non-invasive blood pressure estimation is increasingly recognized for its integral role in enhancing CDSS capabilities [10, 11]. Consequently, accurately identifying and removing erroneous waveforms and SpO2 values from CDSS inputs is crucial. Ensuring the reliability of these inputs is essential for the effective functioning of the CDSS, directly impacting patient outcomes and care efficiency. Building on the foundation of our work in fully-supervised and semi-supervised learning methodologies, we have delved into the domain of PPG artifact detection, focusing on machine learning (ML) applications. A pivotal study in this area [12] investigated the use of machine learning techniques for this purpose. However, challenges arise in scenarios featuring imbalanced classes and limited data availability. In these contexts, Transformer models, despite their advanced attention mechanisms, have exhibited suboptimal performance compared to other methods, such as semi-supervised label propagation and supervised KNN learning. The core issue lies in the Transformer models’ reduced efficacy in smaller datasets. To address these limitations and enhance the Transformer’s applicability in artifact detection, our recent study [13] introduced an innovative approach. We incorporated the Gated Residual Network (GRN) into the Transformer framework, enhancing its performance capabilities significantly. This GRN-Transformer hybrid model not only overcomes the inherent limitations of traditional Transformer models in handling smaller datasets but also outperforms other existing models in artifact detection accuracy and reliability. Despite these advancements in artifact detection, a common limitation persists across recent studies: a heavy reliance on annotated data to train fully supervised machine learning algorithms. This reliance is particularly evident considering that in most cases, only up to 10% of the data is annotated, leaving a vast 90% of the data pool unexploited. Such underutilization of available data presents a significant challenge, especially when dealing with small datasets and imbalanced classes, which are common in the task of binary classification for motion artifact detection in PPG signals. In light of these challenges, this study aims to transcend the constraints of labeled data dependency by harnessing the potential of SSL. This approach is particularly pertinent for two key reasons as follows: i) The vast majority of our data, approximately 90%, remain unannotated, representing a largely untapped resource that could significantly enhance our understanding and detection capabilities, and ii) There is a compelling opportunity to explore how SSL can adapt and perform with limited and unlabeled data, a scenario frequently encountered in clinical settings. By pivoting towards SSL, we aim to leverage the underused unannotated data, potentially revolutionizing how we approach artifact detection and enhancing the robustness of our classification models in these challenging PICU environments.
Figure 1: An end-end process diagram workflow demonstration.
In our quest to validate the efficacy of SSL in PPG artifact classification, as shown in Fig. 1, we explored three distinct SSL mechanisms: masking, contrastive learning, and self-distillation without labels (DINO). Masking involves intentionally obscuring parts of the input data, prompting the model to predict these missing parts, thereby learning robust data representations. Contrastive learning focuses on learning by comparing similar and dissimilar pairs of data points, enhancing the model’s ability to discern subtle differences and similarities within the data. DINO, or self-distillation without labels, leverages a teacher-student network architecture where the ‘student’ model learns to mimic the output of a ‘teacher’ model, which is typically more powerful or pre-trained, even without labeled data. In the workflow Fig. 1, the model first (Step 1) undergoes pre-training with techniques like masking, contrastive learning, and DINO on unannotated data, allowing it to capture intrinsic PPG signal patterns. Following this, the pre-trained model is fine-tuned on a labeled dataset (Step 2), focusing on its ability to accurately classify and detect artifacts. This two-tiered approach leverages the extensive unannotated data for foundational learning and a smaller annotated dataset for precision, equipping the model with robust generalization capabilities for effective artifact detection. Consequently, the experimental results underscored a significant finding: employing these SSL techniques to initially train on unlabeled data, followed by fine-tuning on a smaller proportion (10%) of labeled data, markedly improved the classification performance over models trained exclusively on fully-supervised learning with limited labeled data. Specifically, the Transformer model is robust and effective compared to other backbone neural networks when it is elaborated with SSL. Second, contrastive learning emerged as the most effective among the three SSL techniques tested. However, its success is heavily contingent upon the design of the contrastive loss functions. Therefore, we delved further into investigating various contrastive losses. A notable improvement was the adaptation and enhancement of the InfoNCE loss function, which proved to be particularly effective in augmenting the performance of contrastive models. Lastly, unsupervised learning also helps the Transformer’s performance; however, it can not beat SSL’s efficacy in all data proportion cases.
@@ II Related Works
The accuracy and reliability of pulse oximeters, essential for monitoring blood oxygen levels, are significantly impacted by motion artifacts in PPG signals [14]. Traditional filtering algorithms, while helpful, are not fully effective in eliminating these motion artifacts. As a result, residual motion artifacts often lead to inaccuracies in the measurement of blood oxygen saturation. This underscores the importance of developing more sophisticated methods for detecting and mitigating motion artifacts in PPG signals to ensure the reliability and precision of pulse oximetry readings. In the PICU setting, the necessity for accurate PPG artifact detection becomes even more pronounced [12]. Children, especially those in critical care, are often prone to more frequent and abrupt movements compared to adults. Therefore, enhancing PPG signal processing to accurately identify and compensate for these artifacts is vital for improving patient monitoring and care quality in pediatric intensive care settings.
@@
The integration of ML into PPG analysis has revolutionized a range of clinical applications, addressing diverse healthcare needs with greater precision. A notable example is heart rate estimation, where ML models have shown exceptional accuracy, as highlighted in the studies by Dao et al. [15] and Mehrgardt et al. [16]. This advancement extends beyond basic applications; ML also enables real-time physiological monitoring. It provides critical insights into vital signs such as blood pressure [17], oxygen saturation [11], and respiratory rate [18], thereby enhancing patient care and enabling the early identification of potential health issues, as shown in research by Venema et al. [19] and Alharbi et al. [20]. Furthermore, ML algorithms have been particularly effective in the detection and filtration of motion artifacts in PPG data, a crucial step in ensuring the reliability of continuous monitoring systems, as demonstrated by Nwibor et al. [21]. In summary, the application of ML in PPG analysis has not only broadened the scope of clinical applications but has also significantly increased the efficiency and accuracy of healthcare interventions.
@@
The implementation of conventional ML techniques marked initial advancements in this field. For instance, Support Vector Machine classifiers have demonstrated efficacy in detecting heart rates from PPG signals, employing time-frequency spectral features [15]. This approach represents the foundational phase of machine learning applications in PPG analysis. However, the field has witnessed a significant shift towards more sophisticated methods, particularly with the advent of deep learning algorithms. Studies have increasingly adopted deep learning models like Multilayer Perceptrons and Fully Convolutional Neural Networks, which have shown promising results in artifact detection [22, 23]. These advancements reflect a transition from conventional machine learning techniques to more complex and capable deep learning methods within the supervised learning paradigm, significantly enhancing the capabilities in PPG signal analysis.
@@
Recent advancements in PPG signal analysis underscore a paradigm shift from traditional supervised learning methods to exploring the potential of unsupervised and semi-supervised approaches. Initial studies, such as the one by Maqsood et al. [24], have demonstrated the effectiveness of deep learning algorithms, particularly Bi-LSTM combined with time-domain features, in achieving superior heart rate estimation. This approach primarily relied on supervised learning frameworks. However, our research team’s study [12] marks a significant transition in this domain. We explored a range of ML techniques, encompassing not only conventional machine learning and supervised models like MLP and Transformer but also semi-supervised learning with label propagation. Interestingly, we found that semi-supervised label propagation and the supervised KNN algorithm exhibited better performance than Transformer models, especially in scenarios involving imbalanced classes and limited data. This finding indicates a growing trend toward leveraging semi-supervised methods in PPG artifact detection, highlighting their potential to overcome data scarcity and class imbalance challenges.
@@
In conclusion, the integration of ML in PPG analysis has marked a significant milestone in the field, especially in enhancing artifact detection. This advancement has improved signal reliability and opened new frontiers in clinical applications. However, this domain still requires extensive research and development to fine-tune these methodologies and ensure their effective integration into clinical practices. This ongoing progression is set to redefine the landscape of PPG signal analysis and its consequential impact on patient care, particularly in the PICU at CHUSJ. Additionally, despite the transformative potential of Transformer models and attention mechanisms in ML, as surveyed by Lin et al. [26], their efficacy in small dataset scenarios remains a challenge, as highlighted in studies by our studies [12, 16, 27]. This study, therefore, was driven by the objective of augmenting the SSL capability in effectively managing small datasets and imbalanced classes. Our focus is particularly on the classification task of detecting motion artifacts in PPG signals, a critical aspect of ensuring accurate and reliable patient monitoring.
@@ III-A Clinical PPG Data at CHUSJ
The research protocol conducted for this study received approval from the research ethics board of CHUSJ, University of Montreal under the project number eNIMP:2023-4556. The data collection process within this HRDB captures a wide range of physiological signals. Key among these was the pulse oximeter sensor, which played a crucial role in acquiring PPG signals. This sensor works by emitting light into the skin and measuring the variations in light absorption due to blood flow changes during the cardiac cycle. Additionally, blood pressure signals were recorded through non-invasive methods, offering a comprehensive view of blood pressure dynamics.
The study’s population included children aged from newborn to 18 years who were admitted to CHUSJ between September 2018 and September 2023. The inclusion criteria were based on the availability of essential waveform records such as ECG, PPG, and ABP. Certain exclusion criteria were applied to maintain the integrity and quality of the data. Any data collected after the fourth day of hospitalization were excluded to minimize potential biases. Moreover, the study did not include patients undergoing Extracorporeal Membrane Oxygenation (ECMO) treatment. In cases of multiple readmissions, only data from the initial admission were considered to maintain data independence and avoid confounding factors.
The final cohort comprised 1,573 eligible patients. For each patient, continuous recordings of ECG, PPG, blood pressure via catheter, and cuff blood pressure were made over a 96-hour period. The PPG signals were specifically captured every 5 seconds at a sampling frequency of 128 Hz. Similarly, blood pressure and ECG signals were acquired with a frequency of 512 Hz, also at 5-second intervals. During the data extraction phase, a fixed 30-second window was employed for each PPG signal to facilitate subsequent processing and analysis.
@@ III-B Data Pre-Processing
In our study, data preprocessing plays a crucial role in enhancing the quality of PPG signals and involves four key steps. This process encompasses four main steps: filtering, segmentation, resampling and normalization, and feature extraction. Initially, each signal is refined using a bandpass Butterworth filter with frequencies between 0.5 Hz and 5 Hz, effectively removing baseline wander and high-frequency noise. Subsequently, the signal is segmented into smaller pulses by identifying local minima, enabling precise artifact detection within each pulse. Each pulse is then uniformly resampled to 256 samples, corresponding to a one-second heart cycle, and normalized to ensure a consistent feature scale. Finally, feature extraction focuses on the signal’s temporal characteristics, with samples taken every four milliseconds to capture the complete temporal profile of each pulse. These preprocessing steps are critical for ensuring data quality and reliability for subsequent analysis.
A manual annotation process is employed to establish ground truth for classification assessment. An expert annotates each PPG signal pulse based on its morphology and characteristics. Additionally, a secondary algorithm, acting as a supplementary expert, reannotates all pulses and employs statistical analysis to determine normative pulse values. The comparison of annotations from both the expert and the algorithm enhances the accuracy and reliability of artifact detection, thereby ensuring robust analysis.
In our final phase of refining the machine learning algorithms for automated artifact classification, we conducted experiments with different proportions of the dataset being annotated. This approach aimed to identify the minimum dataset size necessary for effective annotation. We methodically annotated subsets of 2.5%, 5%, 7.5%, and 10% of the entire dataset, examining each to ascertain the most efficient subset size for optimal performance in artifact detection.
III-C Self-Supervised Learning (SSL)

### v2

@@ Abstract
Recent research has revealed that traditional machine learning methods, such as semi-supervised label propagation and K-nearest neighbors, outperform Transformer-based models in artifact detection from photoplethysmogram (PPG) signals, mainly when data is limited. This study addresses the underutilization of abundant unlabeled data by employing self-supervised learning (SSL) to extract latent features from these data, followed by fine-tuning on labeled data. Our experiments demonstrate that SSL significantly enhances the Transformer model’s ability to learn representations, improving its robustness in artifact classification tasks. Among various SSL techniques—including masking, contrastive learning, and DINO (self-distillation with no labels)—contrastive learning exhibited the most stable and superior performance in small PPG datasets. Further, we delve into optimizing contrastive loss functions, which are crucial for contrastive SSL. Inspired by InfoNCE, we introduce a novel contrastive loss function that facilitates smoother training and better convergence, thereby enhancing performance in artifact classification. In summary, this study establishes the efficacy of SSL in leveraging unlabeled data, particularly in enhancing the capabilities of the Transformer model in PPG artifact detection. This approach holds promise for broader applications in PICU environments, where annotated data is often limited.
@@ I Introduction
Recently the Pediatric Intensive Care Unit (PICU) at CHU Sainte-Justine (CHUSJ) has made notable advancements by developing a high-resolution research database (HRDB) [1, 2]. This innovative database directly integrates biomedical signals from various monitoring devices into the electronic patient record, ensuring seamless data incorporation throughout a patient’s PICU stay [3]. The integration of HRDB has significantly enhanced the Clinical Decision Support System (CDSS) at CHUSJ, boosting patient safety and underpinning decision-making with robust evidence [4]. In this context, early and accurate diagnosis of acute respiratory distress syndrome (ARDS) is a pivotal goal of the CDSS at CHUSJ. Oxygen saturation (SpO2) values, critical in ARDS diagnosis, are central to predicting and managing ARDS [5, 6], and play a key role in determining respiratory support strategies [7, 8, 9]. To have a valid SpO2 value, it is necessary to have a good signal, i.e., Photoplesthymography waveform. Moreover, the PPG waveforms can also be used in algorithms to estimate non-invasive blood pressure [10, 11]. Consequently, accurately identifying and removing erroneous waveforms and SpO2 values from CDSS inputs is crucial. Ensuring the reliability of these inputs is essential for the effective functioning of the CDSS, directly impacting patient outcomes and care efficiency.
@@
Building on the foundation of our work in fully-supervised and semi-supervised learning methodologies, we have delved into PPG artifact detection, focusing on machine learning (ML) applications. A pivotal study in this area [12] investigated the use of machine learning techniques for this purpose. However, challenges arise in scenarios featuring imbalanced classes and limited data availability. In these contexts, Transformer models, despite their advanced attention mechanisms, have exhibited suboptimal performance compared to other methods, such as semi-supervised label propagation and supervised KNN learning. The core issue lies in the Transformer models’ reduced efficacy in smaller datasets. To address these limitations and enhance the Transformer’s applicability in artifact detection, our recent study [13] introduced an innovative approach. We incorporated the Gated Residual Network (GRN) into the Transformer framework, enhancing its performance capabilities significantly. This GRN-Transformer hybrid model not only overcomes the inherent limitations of traditional Transformer models in handling smaller datasets but also outperforms other existing models in artifact detection accuracy and reliability. Despite these advancements in artifact detection, a standard limitation persists across recent studies: a heavy reliance on annotated data to train fully supervised machine learning algorithms. This reliance is particularly evident considering that in most cases, only up to 10% of the data is annotated, leaving a vast 90% of the data pool unexploited. Such underutilization of available data presents a significant challenge, especially when dealing with small datasets and imbalanced classes, which are common in binary classification for motion artifact detection in PPG signals. In light of these challenges, this study aims to transcend the constraints of labeled data dependency by harnessing the potential of SSL. This approach is particularly pertinent for two key reasons as follows: i) The vast majority of our data, approximately 90%, remain unannotated, representing a largely untapped resource that could significantly enhance our understanding and detection capabilities, and ii) There is a compelling opportunity to explore how SSL can adapt and perform with limited and unlabeled data, a scenario frequently encountered in clinical settings. By pivoting towards SSL, we aim to leverage the underused unannotated data, potentially revolutionizing how we approach artifact detection and enhancing the robustness of classifiers in these challenging PICU environments.
Figure 1: An end-end process diagram workflow demonstration for the proposed two-stage framework for PPG signal analysis. Step 1 involves pre-training the encoder on unannotated data using different self-supervised learning techniques (masking, contrastive, and DINO) to learn optimal hidden representations. Step 2 fine-tunes the pre-trained encoder with a classifier using annotated data to optimize the classification loss, thereby improving model performance with minimal labeled data.
In exploring SSL’s effectiveness for PPG artifact classification, as depicted in Fig. 1, we implemented three distinct SSL strategies: masking, contrastive learning, and self-distillation without labels (DINO). Masking involves concealing parts of the input data and training the model to reconstruct these hidden segments, enhancing its capability to capture local and global data features. Contrastive learning, on the other hand, optimizes the model by learning to distinguish between similar and dissimilar data points, refining the representation space to capture subtle patterns in the signal effectively. DINO employs a teacher-student architecture, where the student model mimics the teacher’s output, even without labeled data, promoting the extraction of meaningful representations.
The framework is divided into two stages. The first stage focuses on pre-training the encoder network using various SSL techniques, such as masking, contrastive learning, and DINO. This step aims to learn optimal hidden representations, denoted as $z_{i}$, for the unannotated training data $X_{un-annotated}$. This is achieved by minimizing a contrastive loss specific to each SSL technique, which encourages the encoder to generate meaningful embeddings that capture the intrinsic structure of the data. These embeddings $z_{i},z_{j}$, and $z_{k}$ represent different transformations of the same signal or similar signals, which are contrasted against each other to refine the representation. Second stage, the learned hidden representations from the pre-trained encoder $z_{i}$ are transferred to a classifier network. This step involves fine-tuning the encoder along with the classifier using a small set of annotated data $X_{annotated}$. The goal is to optimize the classification loss, which measures the discrepancy between the predicted labels from the classifier and the true annotated labels. By leveraging the pre-trained representations, the framework improves performance on the classification task with minimal labeled data.
The combined approach ensures the effective utilization of unannotated data during the pre-training stage and maximizes the impact of the limited annotated data during the fine-tuning stage. The experimental results revealed that this combination of pre-training on unlabeled data and fine-tuning on just 10% of annotated data substantially enhanced classification performance compared to models trained using traditional supervised learning on limited data. Contrastive learning demonstrated the highest efficacy among the three SSL methods, underscoring the critical role of contrastive loss function design. To optimize this, we refined the standard InfoNCE loss, leading to a more stable training process and better convergence, significantly boosting the model’s artifact detection performance. Although unsupervised learning techniques contributed positively, they could not match the performance gains achieved with SSL, emphasizing the superiority of leveraging structured pre-training strategies for improving model robustness and generalization.
@@ II Related Works
The accuracy and reliability of pulse oximeters, essential for monitoring blood oxygen levels, are significantly impacted by motion artifacts in PPG signals [14]. Traditional filtering algorithms, while helpful, are not fully effective in eliminating these motion artifacts. As a result, residual motion artifacts often lead to inaccuracies in measuring blood oxygen saturation. This underscores the importance of developing more sophisticated methods for detecting and mitigating motion artifacts in PPG signals to ensure the reliability and precision of pulse oximetry readings. In the PICU setting, the necessity for accurate PPG artifact detection becomes even more pronounced [12]. Children, especially those in critical care, are often prone to more frequent and abrupt movements than adults. Therefore, enhancing PPG signal processing to identify and compensate for these artifacts accurately is vital for improving patient monitoring and care quality in pediatric intensive care settings.
@@
The integration of ML into PPG analysis has revolutionized various clinical applications, addressing diverse healthcare needs with greater precision. A notable example is heart rate estimation, where ML models have shown exceptional accuracy, as highlighted in the studies by Dao et al. [15], and Mehrgardt et al. [16], with accuracies greater than 95%. This advancement extends beyond basic applications; ML also enables real-time physiological monitoring by providing up to over 18% improvements in prediction accuracy. For example, it provides critical insights into vital signs such as oxygen saturation [11], blood pressure [17], and respiratory rate [18], thereby enhancing patient care and enabling the early identification of potential health issues, as shown in research by Venema et al. [19] and Alharbi et al. [20], achieving over 80% for precision, and recall for prediction of patient condition. Furthermore, ML algorithms have been particularly effective in detecting and filtrating motion artifacts in PPG data, a crucial step in ensuring the reliability of continuous monitoring systems, as demonstrated by Nwibor et al. [21] with 98.7% accuracy. Consequently, the application of ML in PPG analysis has not only broadened the scope of clinical applications but has also significantly increased the efficiency and accuracy of healthcare interventions.
@@
The implementation of conventional ML techniques marked initial advancements in this field. For instance, Support Vector Machine classifiers have effectively detected heart rates from PPG signals, employing time-frequency spectral features [15], with accuracies greater than 95%. This approach represents the foundational phase of machine learning applications in PPG analysis. However, the field has witnessed a significant shift towards more sophisticated methods, particularly with the advent of deep learning algorithms. Studies have increasingly adopted deep learning models like Multilayer Perceptrons and Fully Convolutional Neural Networks, which have shown promising results in artifact detection [22, 23], accuracy greater than 90%. These advancements reflect a transition from conventional machine learning techniques to more complex and capable deep learning methods within the supervised learning paradigm, significantly enhancing the capabilities in PPG signal analysis.
@@
Recent advancements in PPG signal analysis underscore a paradigm shift from traditional supervised learning methods to exploring the potential of unsupervised and semi-supervised approaches. Initial studies, such as the one by Maqsood et al. [24], have demonstrated the effectiveness of deep learning algorithms, particularly Bi-LSTM combined with time-domain features, in achieving superior heart rate estimation. This approach primarily relied on supervised learning frameworks. However, our research team’s study [12] marks a significant transition in this domain, achieving a precision of 91%, a recall of 90%, and an F1 score of 90%. We explored a range of ML techniques, encompassing not only conventional machine learning and supervised models like MLP and Transformer but also semi-supervised learning with label propagation. Interestingly, we found that semi-supervised label propagation and the supervised KNN algorithm exhibited better performance than Transformer models, especially in scenarios involving imbalanced classes and limited data. This finding indicates a growing trend toward leveraging semi-supervised methods in PPG artifact detection, highlighting their potential to overcome data scarcity and class imbalance challenges.
@@
In conclusion, the integration of ML in PPG analysis has marked a significant milestone in the field, especially in enhancing artifact detection. This advancement has improved signal reliability and opened new frontiers in clinical applications. However, this domain still requires extensive research and development to fine-tune these methodologies and ensure their effective integration into clinical practices. This ongoing progression is set to redefine the landscape of PPG signal analysis and its consequential impact on patient care, particularly in the PICU at CHUSJ. Additionally, despite the transformative potential of Transformer models and attention mechanisms in ML, as surveyed by Lin et al. [26], their efficacy in small dataset scenarios remains a challenge, as highlighted by our recent studies [12, 16, 27]. This study, therefore, was driven by the objective of augmenting the SSL capability in effectively managing small datasets and imbalanced classes. Our focus is particularly on the classification task of detecting motion artifacts in PPG signals, a critical aspect of ensuring accurate and reliable patient monitoring.
@@ III-A Clinical PPG Data at CHUSJ
The research protocol conducted for this study received approval from the research ethics board of CHUSJ, University of Montreal, under the project number eNIMP:2023-4556. The data collection process within this HRDB captures a wide range of physiological signals. Key among these was the pulse oximeter sensor, which played a crucial role in acquiring PPG signals. This sensor works by emitting light into the skin and measuring the variations in light absorption due to blood flow changes during the cardiac cycle.
The study’s population included children aged from newborn to 18 years who were admitted to CHUSJ between September 2018 and September 2023. The inclusion criteria were based on the availability of essential waveform records such as PPG, electrocardiogram (ECG), and arterial brood pressure (ABP). Certain exclusion criteria were applied to maintain the integrity and quality of the data. Any data collected after the fourth day of hospitalization due to prolonged PICU stay were excluded to minimize potential biases. Moreover, the study did not include patients undergoing Extracorporeal Membrane Oxygenation (ECMO) treatment. In cases of multiple readmissions, only data from the initial admission were considered to maintain data independence and avoid confounding factors.
The final cohort comprised 1,571 eligible patients. For each patient, continuous recordings of ECG, PPG, and blood pressure via catheter were made over a 96-hour period. The PPG signals were specifically captured every 5 seconds at a sampling frequency of 128 Hz. Similarly, blood pressure and ECG signals were acquired with a frequency of 512 Hz, also at 5-second intervals. During the data extraction phase, a fixed 30-second window was employed for each PPG signal to facilitate subsequent processing and analysis.
@@ III-B Data Pre-Processing
In our study, data preprocessing plays a crucial role in enhancing the quality of PPG signals, adhering to the methodologies established in our prior studies [12, 13]. The raw PPG signal undergoes a series of preprocessing steps to enhance signal quality, eliminate noise, and prepare it for subsequent analysis stages. This process encompasses four main steps: filtering, segmentation, resampling, and normalization, detailed as follows:
  1. 1.
    Filtering: Each signal segment is processed using a band-pass Butterworth filter with cut-off frequencies set to 0.5 Hz and 5 Hz, corresponding to heart rates ranging from 30 to 300 beats per minute. A forward-backward filtering approach is applied to preserve the phase integrity of the signal. This step helps eliminate baseline drifts and high-frequency noise components.
  2. 2.
    Pulse segmentation: Local minima detection is employed to segment the signal, defining each pulse between consecutive minima. This segmentation strategy isolates individual pulses, making analyzing and detecting artifacts within each segment easier. The dynamic segment size varies based on signal characteristics and the desired analysis objectives.
  3. 3.
    Resampling: Given that a cardiac cycle in children typically spans 0.3 to 1 second, each pulse is resampled to ensure a uniform representation of 256 data points, equivalent to a 1-second cardiac cycle. Missing values are interpolated using a linear interpolation technique, which is selected for its computational efficiency and ability to maintain signal continuity, thereby preserving the integrity of the pulse waveform.
  4. 4.
    Normalization: The resampled data are normalized to have zero mean and unit variance, ensuring that all features in the dataset are on a comparable scale. This step prevents numerical imbalances from disproportionately influencing subsequent modeling stages.
  5. 5.
    Data transformation: Each segmented pulse is represented as a structured vector containing 256 values, corresponding to the uniformly sampled points derived in the resampling step. This standardized vector format facilitates using the PPG pulses in statistical analyses and machine learning applications, ensuring that the data are consistent and manageable for various downstream processing tasks.
III-C Data Annotation
After segmenting the PPG signals, each pulse is presented to a professional clinician at CHUSJ for annotation. The clinician classifies each pulse as either non-artifact or artifact motion. A pulse is considered artifact-free if its morphology - defined by consistent amplitude, width, and shape - matches that of surrounding pulses. Conversely, a pulse is marked as containing artifacts if it deviates significantly in these characteristics compared to its neighboring pulses.
To validate the annotations provided by the expert, an automated re-annotation algorithm, functioning as a secondary reviewer, was developed [12]. This algorithm utilizes statistical features such as skewness, kurtosis, and standard deviation to differentiate artifact-free pulses from those containing motion artifacts. For consistent cardiac cycles, these values remain stable, while deviations indicate the presence of artifacts. Let $X$ denote the pulse sample values, with $\mu$ as the mean and $\sigma$ as the standard deviation. The skewness, kurtosis, and standard deviation are defined as follows:
$\text{Kurt}[X]$ $=\frac{E\left[\left(X-\mu\right)^{4}\right]}{\sigma^{4}}$ (1)
$\text{Skew}[X]$ $=\frac{E\left[\left(X-\mu\right)^{3}\right]}{\sigma^{3}}$ (2)
$\text{std}[X]$ $=\sqrt{E\left[\left(X-\mu\right)^{2}\right]}$ (3)
To detect anomalies, a 95% confidence interval is established, such that a pulse is classified as normal if its statistical values fall within:
$th_{l}$ $=\mu-2\sigma$ (4)
$th_{u}$ $=\mu+2\sigma$ (5)
If these statistics lie outside the defined thresholds, the pulse is marked as containing artifacts. This approach helps ensure consistency and reduces the risk of human error in manual annotations, ultimately improving the reliability of the pulse classification thereby providing robust analysis.
Finally, with an average of 51 pulses per signal and 1,571 signals in the database, approximately 80,000 pulses were available. Previously, only 10% of the pulses were annotated, leaving 90% unannotated [12]. To leverage this unannotated data, we will experiment with varying annotation proportions of 2.5%, 5%, 7.5%, and 10% - to determine the minimal subset required for effective artifact detection. This approach helps identify the optimal dataset size for robust performance.
III-D Self-Supervised Learning (SSL)